# Getting to the Bottom of Serverless Billing

### v1

@@ 2.4. Turnaround Time Billing and Cold Starts
Billing on wall-clock turnaround time has become a common practice to compensate for the initialization phase: The serverless runtime sandbox lifecycle typically consists of initialization (cold start), request execution, keep-alive (KA), and shutdown (e.g., SIGTERM handling) (web, 2025x; Lin and Shahrad, 2024). Depending on whether the initialization duration is included, serverless providers usually define billable wall-clock time as either execution time or turnaround time (i.e., execution time plus initialization) in their request-based, pay-per-use billing models. Besides billing based on execution time and turnaround time, users may customize provisioned concurrency, minimum instances, or scale-down delay, and pay for the whole runtime instance lifespan (i.e., instance time) on most platforms (web, 2025g). Such instance time billing can further increase billable resources under bursty traffic patterns since scale-down-to-zero is delayed or disabled, and instance idle time is billed. We observe that billing based on turnaround time has become increasingly popular on major platforms: GCP and IBM explicitly state that they charge for the turnaround time (Fassnacht and Regge, 2021; web, 2025t), while AWS recently updated its billing model to include the initialization phase starting August 2025 (Gupta and Gebhart, 5 07).
@@
Our analysis of cold start resource usage helps explain why providers favor turnaround time billing. We analyze 388,955 traceable cold starts from the Huawei serverless traces (web, 2024b; Joosen et al., 2025). For each cold start, we consider the duration spent on the initialization of the runtime sandbox and the resource allocation. We compute the difference between the billable resources (measured in wall-clock resource allocations) consumed during cold start and the sum of billable resources used by all subsequent requests within the sandbox. A negative difference means the cold start alone consumed more billable resources than all later requests combined. Figure 4 quantifies, for the first time, such relative resource cost of cold starts compared to subsequent executions on production systems, which shows that 42.1% of cold starts produce a zero or negative difference. In other words, under a billing model based on purely execution duration of requests, providers would charge less (or the same) for request execution than the actual cost of the initialization phase in about 42.1% of cold start cases.
@@ 2.5. High Invocation Fee and Expensive Rounding Up
Figure 5. The equivalent billable wall-clock time of invocation fees (left) and the rounding up amount of billable wall-clock time and memory (right) (as of 2025-05-15).
Major serverless platforms charge a fixed fee per invocation, typically between \$$1.5\times 10^{-7}$ and \$$6\times 10^{-7}$ per request (web, 2025y, f, v). Although these amounts seem small, they can add up disproportionately when functions run for very short durations or use minimal resources. For example, on AWS Lambda, a fixed invocation fee of \$$2\times 10^{-7}$ is equivalent to 96 ms of billable wall-clock time for a function with the default 128 MB memory configuration, which exceeds the average execution durations reported in the Huawei traces (i.e., 58.19 ms) (Joosen et al., 2025). Figure 5-left shows how invocation fees convert to equivalent billable wall-clock time across different platforms. Besides invocation fees, several platforms apply coarse billing granularity (minimum billing increments) or cutoffs.
The charts on the right in Figure 5 present the inflated billable time and memory usage under different billing granularities for $0.527$ billion requests with execution times of at least 1 ms in Huawei traces (Joosen et al., 2025; web, 2024b). For a 100 ms billing granularity (e.g., GCP and IBM), the average rounded‑up wall-clock time is 77.12 ms, while for a 1 ms granularity with a 100 ms minimum cutoff (e.g., Azure Consumption), the average rounded‑up wall-clock time is 61.35 ms. When billing memory with a 128 MB granularity (e.g., Azure Consumption), the average rounded‑up billable memory is $2.67\times 10^{-2}$ GB‑seconds. These values are on the same order of magnitude as the average execution durations and billable memory amounts reported in the studied serverless traces (58.19 ms and $2.75\times 10^{-2}$ GB‑seconds).
Therefore, our analysis shows that invocation fees are high, and together with billing granularity, they can cause disproportionate costs for short, small function invocations. These extra costs may not be explained only as a way to offset resource usage during the initialization phase (cold start), since providers that bill for turnaround time still charge the invocation fee. They may further be linked to overheads in the serving architecture and OS scheduling, which we analyze further in §3 and §4.
@@ 3.1. Cost Implications of the Concurrency Model
Serverless platforms vary in how they map concurrent requests to sandboxes. Depending on the maximum number of concurrent requests allowed per sandbox, there are two common serving models: a concurrency limit of strictly one (single-concurrency model) and a concurrency limit that can be greater than one (multi-concurrency model).
@@
Under the multi‐concurrency model (e.g., GCP, Azure, and IBM), multiple requests can enter and be executed within the same sandbox concurrently, if the user code supports concurrency. Platforms using this model usually allow users to set the maximum concurrency per sandbox and concurrency-based scaling policies. However, if the extra control knob on concurrency is not configured properly, multi-concurrency can degrade function performance while increasing cost since resource contention (e.g., CPU, memory, and cache) slows down all concurrent requests and increases billable wall time (e.g., execution time). For example, running two CPU-bound requests, each requiring 1 s of CPU time, together in a sandbox with 1 vCPU doubles the execution duration of each request to 2 s, thus doubling billable resources as well. In practice, such slowdowns stemming from resource contention are often worse due to context switch overhead and cache misses (Shahrad et al., 2019).
@@
To show how concurrency models affect performance and cost, we deploy the same compute-intensive function (PyAES from Functionbench (Kim and Lee, 2019)) on AWS Lambda and GCP with 1 vCPU allocation. Each request takes about 160 ms of CPU time. We use the default concurrency configurations (i.e., limit of 80) and scaling policy (60% CPU utilization target and concurrency-based scaling (web, 2025a)) on GCP. At various request rates (in RPS), we send bursts of requests for 120 s to simulate a short traffic spike. The plot on the left in Figure 6 presents the average execution duration reported by providers. AWS maintains a stable execution time at all request rates due to dedicated sandboxes without resource contention. The average execution time (and cost) of the function deployed on GCP rises by up to 9.65×9.65$\times$9.65 × when the request rate is higher than 6 RPS.
@@
Longer traffic logs reveal a key caveat of multi-concurrency models: it takes time to gather scaling metrics to scale instances to match demand. We send a steady traffic of 15 RPS to the GCP function for 20 minutes. The plot on the right in Figure 6 shows the first five minutes (later data remains stable) of execution time and container count reported by GCP. Scaling does not begin until about 40 s. This is likely due to the fact that platforms with the multi-concurrency serving model usually aggregate the scaling metrics over a time window (e.g., 60 s by default in Knative (Knative, 5 07)) to avoid oscillation (web, 2025b). The execution duration and instance count remain stable after around 90 s, but the average duration is still 1.43 times higher (i.e., 239.29 ms versus 166.78 ms) than that under the RPS of 1 due to resource contention.
@@
Figure 6. The multi-concurrency serving model can lead to non-linear slowdowns and increased costs under high concurrency, mainly due to delays in scaling the number of sandboxes to match the incoming request load.
@@ 3.2. Request Serving Architecture Overhead
Figure 7 depicts three common request serving architectures on major serverless platforms. In the runtime API long polling model (e.g., AWS Lambda), the user provides a handler method (non-HTTP) or a binary executable for processing requests. A runtime program (usually offered by providers) runs inside the sandbox and repeatedly polls the runtime API endpoint over HTTP or RPC in a blocking infinite loop. The retrieved request event is processed by the handler, and the result is posted back to the runtime API before the next poll. In the HTTP server model (e.g., Azure, GCP, IBM, and Knative), the function itself runs an HTTP server on a given port, with the user logic wrapped in an HTTP handler. The queue (e.g., usually running in a sidecar container) that receives the request from the ingress acts like a reverse proxy and forwards requests to the HTTP server. In the code/binary execution architecture (e.g., Cloudflare Workers), the user uploads a code block or precompiled binary (e.g., Wasm modules (web, 2025z)). For each request, the language runtime engine (e.g., V8 JavaScript engine) compiles and executes (JIT) or loads and executes the binary, captures the output, and sends back the response.
Figure 8. The functions with HTTP servers have the highest overhead, while code/binary execution has the smallest.
@@
To quantify the overhead of these architectures, we deploy a minimal serverless function that simply returns an empty string and status code across major platforms. Figure 8 presents the execution duration of the minimal function reported by the providers, which indicates the latency added by the request serving architecture, such as polling request events, HTTP routing, and sending back responses. Our measurements reveal that platforms using the HTTP server architecture (i.e., GCP and Azure) usually have notably higher overhead, compared to API polling and code/binary execution architectures, with an average latency up to 5.93 ms. The resource configuration may also affect such overhead (i.e., GCP 0.08 vCPUs vs GCP 1 vCPU), as the HTTP request–response cycle and HTTP server involve CPU-bound tasks (e.g., header and payload parsing, encoding, and serialization), and a lower CPU allocation can slow these operations. The AWS Lambda functions that use long polling maintain a stable overhead of around 1.17 ms on average. Cloudflare Workers delivers near-zero latency (falling below the precision limit of 0.01 ms reported by Cloudflare), suggesting the high efficiency of the code/binary execution architecture.
@@ 3.3. Keep-Alive Duration and Resource Allocation Patterns
Figure 9. The probabilities of having cold starts increase as the function sandbox idle time becomes longer. The keep-alive durations vary across platforms (as of 2025-05-15).
@@
Cold start is one of the main causes of performance degradation in serverless functions, and keep-alive (KA) has become a common practice to mitigate the cold start latency (Shahrad et al., 2020b). Major serverless providers keep the user function sandbox active for a period after each request to reduce the chance of cold starts for subsequent invocations. Common keep-alive mechanisms include scale-down delay (e.g., Azure Consumption Plan, GCP, and IBM) (web, 2025a), container snapshotting (Lan et al., 2024), code caching (e.g., Cloudflare Workers) (Partovi, 2020), and runtime freezing (e.g., AWS Lambda) (web, 2025x). Function keep-alive has a direct impact on provider cost, as idle functions can hold active resources (e.g., memory) or reserved capacity (for some schedulers), affecting deployment density. Even techniques that deallocate CPU and memory during the KA phase, such as snapshotting, freezing (e.g., microVM pause), and caching, require CPU time for processing snapshots and cache/storage space. These costs are ultimately passed on to users through per-unit resource pricing or invocation fees.
@@
We deploy serverless functions on major serverless platforms and analyze the keep-alive durations as well as the underlying keep-alive mechanisms. We send requests at different idle intervals to check whether the sandbox is re-created and empirically measure the keep-alive duration. The idle interval is the duration between the end of the previous invocation and the arrival of the next. Figure 9 presents the probability of a cold start as a function of the sandbox idle time, calculated over 100 data points per idle interval. The results show that AWS Lambda keeps the function sandbox alive for up to 300 to 360 s. Azure is likely to use an opportunistic keep-alive strategy, resulting in varying KA durations between 120 s and 360 s. Also, Azure pre-warms the function if the platform detects cold starts occurring at regular intervals (i.e., through idle time histograms) (Shahrad et al., 2020a; Almeida, 2024). However, we did not observe such behavior in our experiments, despite regular traffic patterns, as we encountered consistent cold starts at high idle times. This is probably due to the test period being too short for Azure to learn traffic patterns (Shahrad et al., 2020b). Besides, Azure may further increase the KA duration for functions with higher traffic and that have been scaled up to multiple instances. We observe a maximum KA duration of around 740 s for the Azure function that has scaled up to 3 instances. In contrast, GCP has the longest keep-alive duration, with the most instances being kept alive for about 900 s. Compared with the data reported in 2018 (e.g., AWS Lambda usually kept functions alive for 27 minutes) (Wang et al., 2018), our observations reveal that keep-alive durations on current serverless platforms still vary but have become shorter than previous measurements, possibly reflecting opportunistic strategies or measures for cost savings.
@@
(wait for SIGTERM handling) (web, 2021)
@@
To examine resource consumption during keep-alive, we run CPU profiling workloads (i.e., Algorithm 1 discussed in §4) and empirically measure the CPU resources available to sandboxes during the KA phase. Table 2 summarizes the resource allocation patterns of the function sandbox during KA and its graceful shutdown behavior when exiting the KA phase and being terminated across platforms. The resource allocation behavior during KA varies across platforms, and so do its performance and cost implications for serverless providers and users.
@@
AWS freezes the sandbox (e.g., puts the microVM to sleep) and resumes it when a new request arrives (Khan, 2021). Therefore, no active CPU or memory resources are allocated during KA. On GCP, CPU allocation is dynamically scaled down to around 0.01 vCPUs during KA and is scaled back up to the user-configured level when requests arrive within the KA window. Both AWS and GCP deallocate or scale down resources during keep-alive, which naturally saves cost. In contrast, Azure appears to make no change to CPU and memory allocation during KA, which may explain its shorter, opportunistic KA period that reflects a trade-off between resource use and cold start probabilities. Cloudflare pre-warms functions on receiving the TLS handshake before the connection establishment, which can mask the very short loading and JIT compilation latency (e.g., around 5 ms) in case of cold starts (Partovi, 2020).
@@
The KA duration and behaviors can directly affect function performance and user costs. A longer KA period can help reduce user costs on platforms that charge based on turnaround time, which includes the cold start latency. Also, keeping resources and sandboxes active during KA (e.g., Azure and GCP) can enable the execution of background tasks and increase the probability of reusing long-lived, persistent connections. Deallocating resources (e.g., AWS and Cloudflare) may cause remote servers to close connections when they stop receiving heartbeat packets, adding overhead and cost when connections must be re-established. Furthermore, Azure maintains full resource allocation during KA, enabling user-created background tasks to run outside the request execution window. Since Azure Consumption is billed on memory consumption during request execution, resource consumption during KA is not billed (although background tasks can still affect billable wall time and memory usage during request executions due to resource contention) (web, 2023a). This may provide opportunities for users to exploit resource allocation. For example, a short request can start a background task running in another thread or coroutine. The background task can send results to other cloud services (e.g., block storage) or remote endpoints after completion, or a subsequent request can retrieve those results. We have successfully implemented this execution pattern on Azure. By doing so, only one or two brief requests are billed, which can substantially reduce overall cloud costs.
@@
Figure 10. The difference between the ideal (expected) and actual execution duration shows CPU over-allocation for functions hosted on major serverless platforms. GCP logs show two sets of quantization jumps, which may be the cause of CPU scaling down/up when entering/exiting KA phases (§3.3).

### v2

@@ 2.4. Turnaround Time Billing and Cold Starts
(I4) Billing on wall-clock turnaround time has become a common practice to compensate for the initialization phase: The serverless runtime sandbox lifecycle typically consists of initialization (cold start), request execution, keep-alive, and shutdown (e.g., SIGTERM handling) (Services, 2025h; Lin and Shahrad, 2024). Depending on whether the initialization duration is included, serverless providers usually define billable wall-clock time as either execution time or turnaround time (i.e., execution time plus initialization) in their request-based, pay-per-use billing models. Besides billing based on execution time and turnaround time, users may customize provisioned concurrency, minimum instances, or scale-down delay, and pay for the whole runtime instance lifespan (i.e., instance time) on most platforms. Such instance time billing can further increase billable resources under bursty traffic patterns since scale-down-to-zero is delayed or disabled, and instance idle time is billed. We observe that billing based on turnaround time has become increasingly popular on major platforms: GCP and IBM explicitly state that they charge for the turnaround time (Fassnacht and Regge, 2021; Cloud, 2025l), while AWS recently updated its billing model to include the initialization phase (cold start delay) starting August 2025 (Gupta and Gebhart, 5 07).
@@
Our analysis of cold start resource usage helps explain why providers favor turnaround time billing. We analyze 388,955 traceable cold starts from the Huawei serverless traces (Joosen, 2025; Joosen et al., 2025). For each cold start, we consider the duration spent on the initialization of the runtime sandbox and the resource allocation. We compute the difference between the billable resources (measured in wall-clock resource allocations) consumed during cold start and the sum of billable resources used by all subsequent requests within the sandbox. A negative difference means the cold start alone consumed more billable resources than all later requests combined. Figure 4 quantifies, for the first time, such relative resource cost of cold starts compared to subsequent executions on production systems, which shows that 42.1% of cold starts produce a zero or negative difference. In other words, under a billing model based on purely execution duration of requests, providers would charge less (or the same) for request execution than the actual cost of the initialization phase in about 42.1% of cold start cases.
@@ 2.5. High Invocation Fee and Expensive Rounding Up
Figure 5. The equivalent billable wall-clock time of invocation fees (left) (as of 2025-05-15) and the rounded-up amount of billable wall-clock time and memory (right).
Major serverless platforms charge a fixed fee per invocation, typically between $1.5\times 10^{-7}$ and $6\times 10^{-7}$ per request (Vercel, 2025; Cloud, 2025a; Services, 2025f). Although these amounts seem small, they can add up disproportionately when functions run for very short durations or use minimal resources. For example, on AWS Lambda, a fixed invocation fee of $2\times 10^{-7}$ is equivalent to 96 ms of billable wall-clock time for a function with the default 128 MB memory configuration, which exceeds the average execution durations reported in the Huawei traces (i.e., 58.19 ms) (Joosen et al., 2025). Figure 5-left shows how invocation fees convert to equivalent billable wall-clock time across different platforms. Besides invocation fees, several platforms apply coarse billing granularity (minimum billing increments) or cutoffs.
The charts on the right in Figure 5 present the inflated billable time and memory usage under different billing granularities for $527.05$ million requests with execution times of at least 1 ms in Huawei traces (Joosen et al., 2025; Joosen, 2025). For a 100 ms billing granularity (e.g., GCP and IBM), the average rounded‑up wall-clock time is 77.12 ms, while for a 1 ms granularity with a 100 ms minimum cutoff (e.g., Azure Consumption), the average rounded‑up wall-clock time is 61.35 ms. When billing memory with a 128 MB granularity (e.g., Azure Consumption), the average rounded‑up billable memory is $2.67\times 10^{-2}$ GB‑seconds. These values are on the same order of magnitude as the average execution durations and billable memory amounts reported in the studied serverless traces (58.19 ms and $2.75\times 10^{-2}$ GB‑seconds).
Therefore, our analysis shows that (I5) invocation fees are high, and together with billing granularity, they can cause disproportionate costs for short, small function invocations. These extra costs may not be explained only as a way to offset resource usage during the initialization phase (cold start), since providers that bill for turnaround time still charge the invocation fee. They may further be linked to overheads in the serving architecture and OS scheduling, which we analyze further in §3 and §4.
@@ 3.1. Cost Implications of the Concurrency Model
Serverless platforms vary in how they map concurrent requests to sandboxes. Depending on the maximum number of concurrent requests allowed per sandbox, there are two common serving models: the single-concurrency model, in which the concurrency limit is strictly one (i.e., no intra-runtime concurrency), and the multi-concurrency model, where the concurrency limit can be greater than one.
@@
Under the multi‐concurrency model, multiple requests can enter and be executed within the same sandbox concurrently, if the user code supports concurrency. Platforms using this model usually allow users to set the maximum concurrency per sandbox and concurrency-based scaling policies. However, (I6) if the extra control knob on concurrency is not configured properly, multi-concurrency can degrade function performance while increasing cost since resource contention (e.g., CPU, memory, and cache) slows down all concurrent requests and increases billable wall-clock time (e.g., execution time) under request-based billing. For example, running two CPU-bound requests, each requiring 1 s of CPU time, together in a sandbox with 1 vCPU doubles the execution duration of each request to 2 s, thus doubling billable resources as well. In practice, such slowdowns stemming from resource contention are often worse due to context switch overhead and cache misses (Shahrad et al., 2019).
@@
To show how concurrency models affect performance and cost, we deploy the same compute-intensive function (PyAES from Functionbench (Kim and Lee, 2019)) on AWS Lambda and GCP with 1 vCPU allocation. Each request takes about 160 ms of CPU time. We use the default concurrency configurations (i.e., limit of 80) and scaling policy (60% CPU utilization target and concurrency-based scaling (Cloud, 2025d)) on GCP. At various request rates (in RPS), we send bursts of requests for 120 s to simulate a short traffic spike. The plot on the left in Figure 6 presents the average execution duration reported by providers. AWS maintains a stable execution time at all request rates due to dedicated sandboxes without resource contention. The average execution time (and cost) of the function deployed on GCP rises by up to $9.65\times$ when the request rate is higher than 6 RPS.
@@
Longer traffic logs reveal a key caveat of multi-concurrency models: it takes time to gather scaling metrics to scale instances to match demand. We send a steady traffic of 15 RPS to the GCP function for 20 minutes. The plot on the right in Figure 6 shows the first five minutes (later data remains stable) of execution time and container count reported by GCP. Scaling does not begin until about 40 s. This is likely due to the fact that platforms with the multi-concurrency serving model usually aggregate the scaling metrics over a time window (e.g., 60 s by default in Knative (Knative, 5 07)) to avoid oscillation (Azure, 2025a). The execution duration and instance count remain stable after around 90 s, but the average duration is still 1.43 times higher (i.e., 239.29 ms versus 166.78 ms) than that under the RPS of 1 due to resource contention.
Such a dual penalty of slowdowns and higher bills stemming from the multi-concurrency model is particularly concerning, given that a recent characterization study reports that 93.3% of serverless workloads on IBM Cloud Code Engine used Knative’s default container concurrency of 100 as the limit (Nasiri et al., 2026; Knative, 2025a). This suggests that most users either do not optimize, or might be unaware of concurrency settings, highlighting the critical need for relevant control knob tuning tools and concurrency optimization guidance from cloud providers.
@@
Figure 6. Function execution durations under varying request rates. The multi-concurrency serving model can lead to non-linear slowdowns and increased costs under high concurrency, mainly due to delays in scaling the number of sandboxes to match the incoming request load.
@@ 3.2. Request Serving Architecture Overhead
Figure 7 depicts three common request serving architectures on major serverless platforms. In the runtime API long polling model (e.g., AWS Lambda (Services, 2025i)), the user provides a handler method (non-HTTP) or a binary executable for processing requests. A runtime program (usually offered by providers) runs inside the sandbox and repeatedly polls the runtime API endpoint over HTTP or RPC in a blocking infinite loop. The retrieved request event is processed by the handler, and the result is posted back to the runtime API before the next poll. In the HTTP server model (e.g., Azure, GCP, IBM, and Knative (Azure, 2025b; Cloud, 2025j, n; Knative, 2025b)), the function itself runs an HTTP server on a given port, with the user logic wrapped in an HTTP handler. The queue (e.g., usually running in a sidecar container) that receives the request from the ingress acts like a reverse proxy and forwards requests to the HTTP server. In the code/binary execution architecture (e.g., Cloudflare Workers), the user uploads a code block or precompiled binary (e.g., Wasm modules (Cloudflare, 2025c)). For each request, the language runtime engine (e.g., V8 JavaScript engine) compiles and executes (JIT) or loads and executes the binary, captures the output, and sends back the response.
Figure 8. Execution durations of the minimal serverless function across platforms with different serving architectures. The functions with HTTP servers have the highest overhead, while code/binary execution has the smallest.
@@
To quantify the overhead of these architectures, we deploy a minimal serverless function that simply returns an empty string and status code across major platforms. We measure the execution duration of the minimal function that encapsulates pod-/container-level system software. Figure 8 presents the execution duration of the minimal function reported by the providers, which reflects the latency added by the request serving architecture, such as polling request events, HTTP routing, and sending back responses. Our measurements reveal that (I7) platforms using the HTTP server architecture (i.e., GCP and Azure) usually have notably higher overhead, compared to API polling and code/binary execution architectures, with an average latency up to 5.93 ms. This added latency can impact short functions or round the billable time up to the next interval.
This is due to that functions with the HTTP server model usually host standard HTTP servers as the upstream of the ingress, queue, and/or load balancer, which add overheads, such as maintaining HTTP listeners, connections, thread pools, and handler routing. Also, requests usually traverse additional middleware (e.g., queues, ingress, and/or service mesh sidecars) across containers and veth devices, adding latencies (Cvetković et al., 2023; Zhu et al., 2023; Cloud, 2025e, k). The resource configuration may also affect such overhead (i.e., GCP 0.08 vCPUs vs GCP 1 vCPU), as the HTTP request–response cycle and HTTP server involve CPU-bound tasks (e.g., header and payload parsing, encoding, and serialization), and a lower CPU allocation can slow these operations. The AWS Lambda functions that use long polling maintain a stable overhead of around 1.17 ms on average. Cloudflare Workers delivers near-zero latency (falling below the precision limit of 0.01 ms reported by Cloudflare), suggesting the high efficiency of the code/binary execution architecture.
@@ 3.3. Keep-Alive Duration and Resource Allocation Patterns
Figure 9. Cold start probabilities versus function idle times. The probabilities of having cold starts increase as the function sandbox idle time becomes longer. The keep-alive durations vary across platforms (as of 2025-05-15).
@@
Cold start is one of the main causes of performance degradation in serverless functions, and keep-alive has become a common practice to mitigate the cold start latency (Shahrad et al., 2020b). Major serverless providers keep the user function sandbox active for a period after each request to reduce the chance of cold starts for subsequent invocations. Common keep-alive mechanisms include scale-down delay (e.g., Azure Consumption Plan, GCP, and IBM) (Cloud, 2025d), container snapshotting (Lan et al., 2024), code caching (e.g., Cloudflare Workers) (Partovi, 2020), and runtime freezing (e.g., AWS Lambda) (Services, 2025h). Function keep-alive has a direct impact on provider cost, as idle functions can hold active resources (e.g., memory) or reserved capacity (for some schedulers), affecting deployment density. Even techniques that deallocate CPU and memory during the keep-alive phase, such as snapshotting, freezing (e.g., microVM pause), and caching, require CPU time for processing snapshots and cache/storage space. These costs are ultimately passed on to users through per-unit resource pricing or invocation fees.
@@
We deploy serverless functions on major serverless platforms and analyze the keep-alive durations as well as the underlying keep-alive mechanisms. We send requests at different idle intervals to check whether the sandbox is re-created and empirically measure the keep-alive duration. The idle interval is the duration between the end of the previous invocation and the arrival of the next. Figure 9 presents the probability of a cold start as a function of the sandbox idle time, calculated over 100 data points per idle interval. The results show that AWS Lambda keeps the function sandbox alive for up to 300 to 360 s. Azure is likely to use an opportunistic keep-alive strategy, resulting in varying keep-alive durations between 120 s and 360 s. Also, Azure pre-warms the function if the platform detects cold starts occurring at regular intervals (i.e., through idle time histograms) (Shahrad et al., 2020a; Almeida, 2024). However, we did not observe such behavior in our experiments, despite regular traffic patterns, as we encountered consistent cold starts at high idle times. This is probably due to the test period being too short for Azure to learn traffic patterns (Shahrad et al., 2020b). Besides, Azure may further increase the keep-alive duration for functions with higher traffic and that have been scaled up to multiple instances. We observe a maximum keep-alive duration of around 740 s for the Azure function that has scaled up to 3 instances. In contrast, GCP has the longest keep-alive duration, with the most instances being kept alive for about 900 s. Compared with the data reported in 2018 (e.g., AWS Lambda usually kept functions alive for up to 27 minutes) (Wang et al., 2018), our observations reveal that (I8) keep-alive durations on current serverless platforms still vary but have become shorter than previous measurements, possibly reflecting opportunistic strategies or measures for cost savings.
@@
(wait for SIGTERM handling) (Services, 2025a)
@@
To examine resource consumption during keep-alive, we run CPU profiling workloads (i.e., Algorithm 1 discussed in §4) and empirically measure the CPU resources available to sandboxes during the keep-alive phase. Table 2 summarizes the resource allocation patterns of the function sandbox during keep-alive and its graceful shutdown behavior when exiting the keep-alive phase and being terminated across platforms. (I9) The resource allocation behavior during keep-alive varies across platforms, and so do its performance and cost implications for serverless providers and users.
@@
AWS freezes the sandbox (e.g., puts the microVM to sleep) and resumes it when a new request arrives (Khan, 2021). Therefore, no active CPU or memory resources are allocated during keep-alive. On GCP, CPU allocation is dynamically scaled down to around 0.01 vCPUs during keep-alive and is scaled back up to the user-configured level when requests arrive within the keep-alive window. Both AWS and GCP deallocate or scale down resources during keep-alive, which naturally saves cost. In contrast, Azure appears to make no change to CPU and memory allocation during keep-alive, which may explain its shorter, opportunistic keep-alive period that reflects a trade-off between resource use and cold start probabilities. Cloudflare pre-warms functions on receiving the TLS handshake before the connection establishment, which can mask the very short loading and JIT compilation latency (e.g., around 5 ms) in case of cold starts (Partovi, 2020).
@@
The keep-alive duration and behaviors can directly affect function performance and user costs. A longer keep-alive period can help reduce user costs on platforms that charge based on turnaround time, which includes the cold start latency. Also, keeping resources and sandboxes active during keep-alive (e.g., Azure and GCP) can enable the execution of background tasks and increase the probability of reusing long-lived, persistent connections. Deallocating resources (e.g., AWS and Cloudflare) may cause remote servers to close connections when they stop receiving heartbeat packets, adding overhead and cost when connections must be re-established. Furthermore, Azure maintains full resource allocation during keep-alive, enabling user-created background tasks to run outside the request execution window. Since Azure Consumption is billed on memory consumption during request execution, resource consumption during keep-alive is not billed (although background tasks can still affect billable wall-clock time and memory usage during request executions due to resource contention) (Ziegler and Henderson, 2023). This may provide opportunities for users to exploit resource allocation. For example, a short request can start a background task running in another thread or coroutine. The background task can send results to other cloud services (e.g., block storage) or remote endpoints after completion, or a subsequent request can retrieve those results. We have successfully implemented this execution pattern on Azure. By doing so, only one or two brief requests are billed, which can substantially reduce overall cloud costs.
@@
Figure 10. Function execution durations and varying fractional CPU allocations. The difference between the ideal (expected) and actual execution duration shows CPU overallocation for functions hosted on major serverless platforms. GCP logs show two sets of quantization jumps, which may be the cause of CPU scaling down/up when entering/exiting keep-alive phases (§3.3).